OpenStudy (anonymous):

At a certain dinner party, ten guests are to be seated around a circular table. How many different arrangements are possidble? (Two arrangements are considered the same if each guest is seated between the same people.)

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

the circle one has always thrown me, but i believe the adjustment is similar to combination.

OpenStudy (amistre64):

10 ways for the first seat, 9 for the next, 8,7,6,5,4,3,2,1 and each arrangement can be resat 10 times due to a shift in seating

OpenStudy (amistre64):



OpenStudy (amistre64):

(1,2) = (2,1) 2!/2 = 1! ------------------------ (1,2,3) = (2,3,1) = (3,1,2) (1,3,2) = (2,1,3) = (3,2,1) 3!/3 = 2! ------------------------- (1,2,3,4) = (2,3,4,1) = (3,4,1,2) = (4,1,2,3) (1,2,4,3) = (2,4,3,1) = (3,1,2,4) = (4,3,1,2) (1,3,4,2) = (2,1,3,4) = (3,4,2,1) = (4,2,1,3) (1,3,2,4) = (2,4,1,3) = (3,2,4,1) = (4,1,3,2) (1,4,2,3) = (2,3,1,4) = (3,1,4,2) = (4,2,3,1) (1,4,3,2) = (2,1,4,3) = (3,2,1,4) = (4,3,2,1) 4!/4 = 3! seems to suggest a pattern of events eh